# Bulk-Boundary Correspondence in 2D Topological Photonics: Analysis and Simulation

### v2

@@ Abstract
The centerpiece of topological photonics is the bulk-boundary correspondence principle (BBCP), which relates discrete invariants of the Bloch bands to the possible presence of topologically protected interface modes between two periodic heterostructures. In addition to the fundamental significance of BBCP, interface modes are of interest in a variety of applications. In Maxwell’s electrodynamics, the BBCP has been rigorously proved for 1D problems, but the 2D case is qualitatively different, because the boundary conditions involve nontrivial Dirichlet-to-Neumann maps rather than scalar impedances as in 1D. The theoretical analysis and numerical examples presented in the paper are consistent with the BBCP. Moreover, the BBCP is closely connected with the positivity of electromagnetic energy density, as has also been shown to be the case in 1D and 1.5D.
@@ I Introduction
If two different periodic structures abut, there may or may not exist modes, either standing or traveling along the interface boundary and confined to that interface. The centerpiece of topological photonics is the bulk boundary correspondence principle (BBCP), which relates the possible presence of interface modes in a bandgap to the discrete invariants of all Bloch bands below that gap. In addition to its fundamental significance, BBCP has practical applications, since topologically protected modes are of interest in communication, topological lasers, and other areas [2, 6, 7, 8].
@@
Furthermore, the BBCP relates evanescent modes within a particular bandgap [^1] More precisely, within a frequency overlap of the bandgaps of the two adjacent heterostructures. to propagating waves at all lower frequencies. Indeed, the relevant parameter is the gap Chern number or gap Zak phase – that is, the sum of Chern numbers (Zak phases) over all bands below a given gap. For example, the existence of evanescent boundary modes at optical frequencies depends, in some way, on propagating waves at radio frequencies, power frequencies, and even on the zero-frequency (static) field. This is perplexing and needs to be understood from both mathematical and physical perspectives.
@@
The present paper deals exclusively with continuous Maxwell electrodynamics. Over the last ten years, the BBCP was rigorously proved for lossless media in 1D and 1.5D [19, 20, 21, 22]. “1.5D” refers to the case where material parameters depend only on the coordinate $n$ normal to the interface; mathematically, the problem can be reduced to one dimension, but the fields and wave vectors may have two components. A central role in the analysis and proof of the BBCP in 1D is played by the Bloch boundary impedance or, alternatively, admittance (the inverses of one another) [19, 20, 21, 22]. In 2D, the scalar admittance turns, mathematically, into an infinite dimensional (in the functional-analytic sense) Dirichlet-to-Neumann (DtN) map (Sec. III), which can be expressed, e.g., in the basis of spatial harmonics along the interface.
@@ II Formulation of four related problems
A canonical setup is shown in Fig. 1. In this paper, only rectangular lattices are considered, with cells scaled to $a\times a$ squares for convenience. Each of the two periodic heterostructures fills its respective half-space. A fundamental question is this: Under what conditions do there exist one or more modes evanescent in the normal direction $n$ – either standing waves or traveling in the tangential direction $\tau$ along the boundary? Purportedly, the BBCP answers this question; however, as noted in the Introduction, a rigorous proof exists in some cases but, to the best of my knowledge and understanding, not for 2D Maxwell problems.
Figure 1: A schematic illustration of two abutting structures; $s$ mode is assumed. The purple vertical arrows indicate a possible direction of propagation of a boundary mode, if it exists.
The main part of the paper focuses on the purest case of topological electrodynamics, whereby losses and frequency dependence (“frequency dispersion”) of material parameters are neglected. The lossy case falls under the umbrella of non-Hermitian photonics [5] and, in 1D, is the subject of Ref. [23]. Accounting for frequency dispersion in my analysis is not conceptually difficult, as demonstrated by the derivation in the Appendix; the relevant techniques have already been worked out in 1D [22] and in 1.5D [24].
@@
as long as both points $(n+m_{n}a,\tau+m_{\tau}a)$ and $(n,\,\tau)$ lie within that structure. For brevity, we assume that the medium right next to the interface boundary is nonmagnetic: $\rho(n=0,\tau)$ = 1. The lattice cells are assumed to be mirror-symmetric with respect to $n$:
@@
and the same for $\mu$. The overall setup is schematically represented in Fig. 1. The boundary modes satisfy Maxwell’s equations (1), (2), with Bloch boundary conditions (8) on the $\tau$ sides of a single semi-infinite strip $S=[0,\infty]\times[-a/2,a/2]$, and decaying to zero in the normal direction on both sides.
Four interrelated auxiliary problems play an important role in the analysis. One is the Bloch problem over a single lattice cell $\mathrm{\Omega}_{C}:0\leq n\leq a;~{}-a/2\leq\tau\leq a/2$:
$e(n+a,\,\tau)=\exp(i\beta a)\,e(n,\,\tau);\quad\quad\partial_{n}e(n+a,\,\tau)=\exp(iqa)\,\partial_{n}e(n,\,\tau)$ (7)
@@
The second and third auxiliary problems (Sec. III) have to do with the DtN operators on the boundary of a single lattice cell and a semi-infinite strip $S=[0,\infty)\times[-a/2,a/2]$, respectively; the strip is composed of an infinite number of cells. Formally, the boundary conditions for $S$ are (8) on the $\tau=\pm a/2$ sides and
@@
The parity of Bloch modes at the $\Gamma$ and $X$ points is known to have major implications in 1D, through a direct connection with the nulls and poles of the Bloch boundary impedance (see [21] for details and references). It is proved in [20, 21, 22] that impedance monotonically decreases as a function of frequency within any band gap.
In 2D, the situation is qualitatively more involved. Admittance turns into a DtN map, defined below and well known in the theory of partial differential equations and in inverse problems. As shown in the Appendix, the DtN operator is monotone as a function of frequency. This generalization of the monotonicity of 1D impedance is interesting but unfortunately not sufficient to prove the BBCP in 2D mathematically.
@@ III Dirichlet-to-Neumann (DtN) operators: Definition and computation
The main stage of the analysis has to do with the behavior of fields in a single semi-infinite strip $S=(0\leq n<\infty;|\tau|\leq a/2)$ composed of repeated lattice cells $C=[0,a]\times[-a/2,a/2]$. The electric field satisfies the generalized Helmholtz equation (2), the Bloch boundary condition at $\tau=\pm a/2$, and the decay condition (10) at infinity. The Dirichlet-to-Neumann (DtN) map $\mathcal{D}$ relates the values of $e$ and $\partial_{n}e$ at the boundary $n=0$ of $S$:
$\partial_{n}e(0,\,\tau;k,\beta)~{}=~{}\mathcal{D}(k,\beta)\,e(0,\,\tau)$ (13)
Since the $n$-derivative does not affect the complex exponential $\exp(i\beta\tau)$, one can easily alternate between $e$ and $\tilde{e}$ (9) in the expressions below. To simplify these expressions slightly, and without much loss of generality, we shall assume that the material immediately adjacent to the interface $n=0$ is nonmagnetic; otherwise the magnetic permeability at $n=0$ would appear in the matching conditions. Function arguments are omitted whenever this should not cause confusion.
@@
$e_{r}=\lambda_{B}e_{l};\quad\partial_{n}e_{r}=\lambda_{B}\,\partial_{n}e_{l}$ (16)
@@
$\left[\lambda_{B}^{2}\mathcal{D}_{lr}^{C}+\lambda_{B}(\mathcal{D}_{ll}^{C}-\mathcal{D}_{rr}^{C})-\mathcal{D}_{rl}^{C})\right]e_{l}~{}=~{}0$ (17)
In [25, 26], analytical solutions for cylindrical inclusions and collocation on the lattice cell boundary are employed. In the present paper, the finite element method (FEM) is used to solve $2(2M+1)$ Dirichlet problems (14) in the Fourier basis $\{\tilde{\psi}_{m}(\tau)\}$ limited to order $m=-M,\ldots,M$. One then obtains a set of $12M+$ 1 Bloch modes – propagating for $|\lambda_{B,m}|=1$ and exponential otherwise. The decaying eigenmodes are then used to construct the DtN operator for the semi-infinite strip $S$. This extension of the DtN-Bloch method [25, 26] was previously proposed and applied to bound states in the continuum at the surface of a photonic crystal in Ref. [27], where further numerical details pertaining to this idea can be found.
The DtN operator can be converted, by a linear transformation, from this evanescent basis to the Fourier basis $\{\tilde{\psi}_{m}(\tau)\}$ along the boundary $n=0$, $-a/2\leq\tau\leq a/2$. The numerical details are not central in the present paper and are omitted. But two main advantages of the FEM-based approach are worth noting: (a) applicability to any composition of the lattice cell, and (b) computational efficiency [^2] The computational cost for the eigenproblem (17) is in practice negligible, since the field at the interface is usually smooth and good accuracy can be attained with a small number of Fourier harmonics. At the same time, sparsity of FE matrices is easier to exploit for Dirichlet solvers than for eigensolvers with large matrices..
With the above definitions in place, we can now turn to the boundary conditions at the interface $n=0$ between the two heterostructures. In 1D, these conditions simply amount to impedance (or admittance) matching. A 2D generalization of these conditions is
@@
Our convention is to treat the two structures on an equal footing, the normal directions in both cases being into the respective domain – hence the plus sign in the expression. The null space, if not empty, contains the boundary value (or values) $e(0,\tau)$ of the possible mode(s) in a chosen basis.
@@ IV Numerical examples
In the simulations reported below, the cell size $a$ of both abutting structures has for convenience been normalized to unity, and the parameters were adjusted to ensure an overlap between the bandgaps of the two heterostructures under investigation. Namely, the permittivity of the gyromagnetic cylinders in the first periodic structure is $\epsilon_{\text{cyl}}=20$, the host medium has $\epsilon_{\text{host}}=2$, and the permeability tensor $\mu$ has, as in [2, 6, 29], $\mu_{nn}=\mu_{\tau\tau}=14$ on the diagonal and $\mu_{n\tau}=-\mu_{\tau n}=12.4i$ off the diagonal. The lattice cell of the second structure contains two rectangular bars $0.2a\times 0.4a$ separated by a $0.2a$ gap (Fig. 2, inset), in an otherwise empty cell. The physical parameters of the bars are $\epsilon_{\text{bar}}=6$, $\mu_{\text{bar}}=6$. Then there exists an overlap in the band gaps of the two structures around the normalized frequency $a/\lambda\sim 0.42$; see the band diagrams in Fig. 2.
@@
The simulations were carried out using FE analysis with first order triangular elements; the typical number of elements was on the order of 5,000, although multiple simulations with finer and coarser meshes were also performed to verify convergence. The DtN map for each of the two structures was computed in the basis of Fourier harmonics as described in Sec. III.
@@
In light of the boundary condition (18), of principal interest is the minimum singular value $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$. Fig. 3 demonstrates a sharp drop in $\sigma_{\min}$ at a particular frequency ($a/\lambda\sim 0.42$) within the overlap between the two band gaps. This sharp drop can be contrasted with the smooth variation of the minimum singular values of the two DtN operators separately (Fig. 3), which indicates that a qualitative change occurs when two structures are put together and share a common interface. The maximum order of spatial harmonics is $M=4$, so that the total number of harmonics between $-M$ and $M$ is nine. Very similar results were obtained with $M=3$ and $M=5$. The mode corresponding to $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$ for $\beta=0$ is illustrated in Fig. 3, right.
Figure 3: Left: The minimum singular value $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$ exhibits a sharp drop at a particular frequency within a gap, in contrast with the smooth variation of the individual minimum singular values of the two DtN operators. Right: the real part of the electric field of the mode corresponding to $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$. The simple and gyromagnetic media occupy the half-spaces $n\leq 0$ and $n\geq 0$, respectively. $\beta=0$; other parameters are specified in the text.
The visual appearance of the modes for the tangential wavenumber $\beta=0$ and $\beta\neq 0$ is similar, as long as a bandgap overlap exists. As an example, cf. surface plots of the real part of the electric field in Fig. 3, right ($\beta=0$) and Fig. 4, left ($\beta=\pi/6$).
@@
The dispersion curve for the interface mode, shown in Fig. 4, right, illustrates an underappreciated point. The difference of the gap Chern numbers $|C_{1}-C_{2}|$ of the respective structures is equal not to the total number of boundary modes but, rather, to the “net current” (a.k.a. “spectral flow” in electronic structure theory) – the difference between the number of modes traveling in one direction and the other [30, p. 5], [31]. The Chern numbers of the gyromagnetic structure are calculated as in [32, 29] and happen to be zero for the first two bands. Hence the Chern number for the second gap (around $a/\lambda\sim 0.42$), which by definition is equal to the sum of the two underlying band Chern numbers, is also zero. The dispersion curve of Fig. 4 indicates the presence of two modes traveling in the opposite directions, implying that the “net flow” at the interface is zero; this is consistent with the gap Chern number being zero.
@@
Figure 4: Left: same as in Fig. 3, right, but for $\beta=\pi/6$. Right: the dispersion curve for the evanescent edge mode. Parameters are specified in the text.

### v3

@@ Abstract
The centerpiece of topological photonics is the bulk-boundary correspondence principle (BBCP), which relates discrete invariants of the Bloch bands to the possible presence of interface modes between two periodic heterostructures. In addition to the fundamental significance of BBCP, interface modes per se are of interest in a variety of applications. In Maxwell’s electrodynamics, the BBCP has been rigorously proved for 1D problems, but the 2D case is qualitatively different, as the boundary conditions involve nontrivial Dirichlet-to-Neumann maps rather than scalar impedances as in 1D. The theoretical analysis and numerical examples in the paper are consistent with the BBCP. Moreover, the BBCP is closely connected with the positivity of electromagnetic energy density, as has also been shown to be true in 1D cases.
@@ I Introduction
If two different periodic structures abut, there may or may not exist modes confined to the interface. The centerpiece of topological photonics is the bulk boundary correspondence principle (BBCP), which relates the possible presence of interface modes in a bandgap to the discrete invariants of Bloch bands. In addition to its fundamental significance, BBCP has practical applications, since topologically protected modes are of interest in communication, topological lasers, and other areas [2, 6, 7, 8].
@@
Furthermore, the BBCP relates evanescent modes within a particular bandgap [^1] More precisely, within a frequency overlap of the bandgaps of the two adjacent heterostructures. to propagating waves at all lower frequencies. Indeed, the relevant parameter is the gap Chern number or gap Zak phase – that is, the sum of Chern numbers (Zak phases) over all bands below a given gap. Somehow, the existence of evanescent boundary modes at, e.g., optical frequencies depends on propagating waves at radio frequencies, power frequencies, and even on the static field. This is perplexing and needs to be understood from both mathematical and physical perspectives.
@@
The present paper deals exclusively with continuous Maxwell electrodynamics. Over the last ten years, the BBCP was rigorously proved for lossless media in 1D and 1.5D [19, 20, 21, 22]. “1.5D” refers to the case where material parameters depend only on the coordinate $n$ normal to the interface; mathematically, the problem can be reduced to one dimension, but the fields and wave vectors may have two components. A central role in the analysis and proof of the BBCP in 1D is played by the Bloch boundary impedance or, alternatively, admittance (the inverses of one another) [19, 20, 21, 22]. In 2D, the scalar admittance turns, mathematically, into an infinite dimensional (in the functional-analytic sense) Dirichlet-to-Neumann (DtN) map (Sec. III).
@@ II Formulation of four related problems
A canonical setup is shown in Fig. 1. In this paper, only rectangular lattices are considered, with cells scaled to $a\times a$ squares for convenience. Each of the two periodic heterostructures fills its respective half-space. A fundamental question is this: Under what conditions do there exist one or more modes evanescent in the normal direction $n$ – either standing waves or traveling in the tangential direction $\tau$ along the boundary? Purportedly, the BBCP answers this question; however, as noted in the Introduction, a rigorous proof exists in some cases but not, to the best of my knowledge and understanding, for 2D Maxwell problems.
Figure 1: An illustration of two abutting structures; $s$ mode is assumed. The purple vertical arrows indicate a possible direction of propagation of an interface mode.
The main part of the paper focuses on the purest case of topological electrodynamics, whereby losses and frequency dependence (“frequency dispersion”) of material parameters are neglected. The lossy case falls under the umbrella of non-Hermitian photonics [5] and, in 1D, is the subject of Ref. [23]. Accounting for frequency dispersion in my analysis is not conceptually difficult (see Appendix); the relevant techniques have already been worked out in 1D [22] and in 1.5D [24].
@@
as long as both points $(n+m_{n}a,\tau+m_{\tau}a)$ and $(n,\,\tau)$ lie within that structure. For brevity, and without much loss of generality, we assume that the medium right next to the interface boundary is nonmagnetic: $\rho(n=0,\tau)$ = 1. The lattice cells are assumed to be mirror-symmetric with respect to $n$:
@@
and the same for $\mu$. The overall setup is schematically represented in Fig. 1. The boundary modes satisfy Maxwell’s equations (1), (2), with Bloch boundary conditions (8) on the $\tau$ sides of a single semi-infinite strip $S=[0,\infty]\times[-a/2,a/2]$, and decaying to zero in the normal direction away from the interface in both structures.
Four interrelated problems need to be considered. One is the Bloch problem over a single lattice cell $C:0\leq n\leq a;~{}-a/2\leq\tau\leq a/2$:
$e(n+a,\,\tau)=\exp(iqa)\,e(n,\,\tau);\quad\quad\partial_{n}e(n+a,\,\tau)=\exp(iqa)\,\partial_{n}e(n,\,\tau)$ (7)
@@
The second and third problems (Sec. III) have to do with the DtN operators on the boundary of a single lattice cell and a semi-infinite strip $S$, respectively; the strip is composed of an infinite number of cells. Formally, the boundary conditions for $S$ are (8) on the $\tau=\pm a/2$ sides and
@@
It is proved in [20, 21, 22] that Bloch impedance monotonically decreases as a function of frequency within any band gap. In 2D, the situation is qualitatively more involved. Admittance turns into a DtN map, defined below and well known in the theory of partial differential equations and in inverse problems. As shown in the Appendix, the DtN operator is monotone as a function of frequency. This generalization of the monotonicity of 1D impedance is interesting but unfortunately not sufficient to prove the BBCP in 2D mathematically.
@@ III Dirichlet-to-Neumann (DtN) operators: Definition and computation
The main stage of the analysis has to do with the behavior of fields in a single semi-infinite strip $S$ composed of repeated lattice cells $C=[0,a]\times[-a/2,a/2]$. The electric field satisfies the generalized Helmholtz equation (2), the Bloch boundary condition at $\tau=\pm a/2$, and the decay condition (10) at infinity. The Dirichlet-to-Neumann (DtN) map $\mathcal{D}$ relates the values of $e$ and $\partial_{n}e$ at the boundary $n=0$ of $S$:
$\partial_{n}e(0,\,\tau;k,\beta)~{}=~{}\mathcal{D}(k,\beta)\,e(0,\,\tau;k,\beta)$ (13)
Since the $n$-derivative does not affect the complex exponential $\exp(i\beta\tau)$, one can easily alternate between $e$ and $\tilde{e}$ (9) in the expressions below. Recall that the material immediately adjacent to the interface $n=0$ is assumed nonmagnetic; otherwise the magnetic tensor at $n=0$ would appear in the matching conditions. Function arguments will be omitted when there is no risk of confusion.
@@
$e_{r}=\lambda_{B}e_{l};\quad\partial_{n}e_{r}=\lambda_{B}\,\partial_{n}e_{l};\quad\quad\lambda_{B}=\exp(iqa)$ (16)
@@
$\left[\lambda_{B}^{2}\mathcal{D}_{lr}^{C}+\lambda_{B}(\mathcal{D}_{ll}^{C}-\mathcal{D}_{rr}^{C})-\mathcal{D}_{rl}^{C})\right]e_{l}~{}=~{}0;$ (17)
In [25, 26], analytical solutions for cylindrical inclusions and collocation on the lattice cell boundary are employed. In the present paper, the finite element method (FEM) is used to solve $2(2M+1)$ Dirichlet problems (14) in the Fourier basis $\{\tilde{\psi}_{m}(\tau)\}$ limited to orders $m=-M,\ldots,M$. One then obtains a set of $2(2M+1)$ Bloch modes – propagating for $|\lambda_{B,m}|=1$ and exponential otherwise. The decaying eigenmodes are then used to construct the DtN operator for the semi-infinite strip $S$. This extension of the DtN-Bloch method [25, 26] was previously proposed and applied to bound states in the continuum at the surface of a photonic crystal in Ref. [27], where further details pertaining to this idea can be found. The DtN operator is then converted, by a linear transformation, from this evanescent basis to the Fourier basis $\{\tilde{\psi}_{m}(\tau)\}$ along the boundary $n=0$, $-a/2\leq\tau\leq a/2$. The numerical procedure is not central to the present paper and is not elaborated upon.
In 1D, Maxwell’s interface conditions between two heterostructures simply amount to impedance (or admittance) matching. A 2D generalization of these conditions is
@@
The plus sign in the expression is due to the fact that the DtN operator in each structure is defined relative to its own normal direction, into the respective domain. The null space, if not empty, contains the boundary value (or values) $e(0,\tau)$ of the possible mode(s) in a chosen basis.
@@ IV Numerical examples
In the simulations reported below, the cell size $a$ of both abutting structures has for convenience been normalized to unity, and the parameters were adjusted to ensure an overlap between the bandgaps of the two heterostructures under investigation. Namely, the permittivity of the gyromagnetic cylinders in the first periodic structure is $\epsilon_{\text{cyl}}=20$, their radius is $r_{\text{cyl}}=0.11a$. The host medium has $\epsilon_{\text{host}}=2$, and the permeability tensor $\mu$ has, as in [2, 6, 28], $\mu_{nn}=\mu_{\tau\tau}=14$ on the diagonal and $\mu_{n\tau}=-\mu_{\tau n}=12.4i$ off the diagonal. The lattice cell of the second structure contains two rectangular bars $0.2a\times 0.4a$ separated by a $0.2a$ gap (Fig. 2, inset), in an otherwise empty cell. The physical parameters of the bars are $\epsilon_{\text{bar}}=6$, $\mu_{\text{bar}}=6$. Then there exists an overlap in the band gaps of the two structures around the normalized frequency $a/\lambda=\omega a/(2\pi c)\sim 0.42$; see the band diagrams in Fig. 2.
@@
The simulations were carried out using FEM with first order triangular elements; the typical number of elements in the lattice cell of each heterostructure was on the order of 5,000, although multiple simulations with finer and coarser meshes were also performed to verify convergence. The DtN map for each of the two structures was computed in the basis of Fourier harmonics as described in Sec. III. Two main advantages of FEM-DtN are worth noting: (a) applicability to any composition of the lattice cell, and (b) computational efficiency [^2] The computational cost for the eigenproblem (17) is in practice negligible, since the field at the interface is usually smooth and good accuracy can be attained with a small number of Fourier harmonics. At the same time, sparsity of FE matrices is easier to exploit in Dirichlet solvers than in eigensolvers with large matrices..
@@
In light of the boundary condition (18), of principal interest is the minimum singular value $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$. Fig. 3 (left) demonstrates a sharp drop in $\sigma_{\min}$ at a particular frequency ($a/\lambda\sim 0.42$) within the overlap between the two band gaps – in contrast with the smooth variation of the minimum singular values of the two DtN operators separately. Thus, a qualitative change occurs when two structures are put together and share a common interface. The maximum order of spatial harmonics is $M=4$, so that their total number is $92M+1=$ 9. Very similar results were obtained with $M=3$ and $M=5$.
@@
The dispersion curve of the interface mode, shown in Fig. 3, right, highlights an often-overlooked point: the difference in gap Chern numbers $|C_{1}-C_{2}|$ of the two structures is equal not to the total number of boundary modes but, rather, to the “net current” (a.k.a. “spectral flow” in electronic structure theory) – the difference between the number of modes traveling in opposite directions [30, p. 5], [31]. The dispersion curve shows two such modes, resulting in a zero net flow; this aligns with the gap Chern number being zero. Indeed, the Chern numbers for the first two bands of the gyromagnetic structure, calculated using the Fukui method [32, 28], are zero. Therefore, the Chern number for the second gap (around $a/\lambda\sim 0.42$) — defined as the sum of the Chern numbers of the underlying bands — is also zero.
Figure 3: Left: The minimum singular value $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$ exhibits a sharp drop at a particular frequency within a gap, in contrast with the smooth variation of the individual minimum singular values of the two DtN operators. Right: the dispersion curve for the evanescent edge mode. Parameters are specified in the text.
Figure 4: The real part of the evanescent electric field corresponding to $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$ for $\beta=0$; other parameters are specified in the text. Left: a surface plot for two lattice cells adjacent to the interface. Right: a color plot for multiple cells; a rapid decay of the field away from the interface is manifest.
The real part of the electric fields corresponding to $\sigma_{\min}(\mathcal{D}_{1}+\mathcal{D}_{2})$ for $\beta a=0$ and $\beta a=\pi/6$ are illustrated in Figs. 4 and 5, respectively. The first field is $\tau$-periodic, while the second one is a wave traveling along the interface.